The base of a certain solid is the region enclosed by the semicircle x^2+y^2=9 for x is less than or equal to 0 and greater than or equal to 3. Every cross section taken perpendicular to the x axis is a semicircle with its diameter across the base. Find the volume of the solid.
1
Expert's answer
2020-09-03T18:24:07-0400
the section of the ball is a circle 

then the solid is the fourth part of the ball

x2+y2+z2=9 R=(9)1/2=3


V=4/3"\pi"R3 /4 = "\pi"R3/3="\pi"(3)3/3= 9"\pi"